An amusement park sells adults'and children's passes.An adult's pass is $25 and a child's pass is $15. A group spent $440 on 20 passes. How many children's passes did the group purchase.
Answer by Boreal(15235)

A+C=20
25A+15C=440
multiply the top by -15
-15A-15C=-300
10A=140
A=14
Therefore, C=6
14*$25=$350
6*$15=$90
$440, checks.
14 adults, 6 children